2 men and 4 boys can do a piece of work in 10 days,
while 4 men and 5 boys can do it in 6 days. Men and
boys are paid wages according to their output. If the
daily wage of a man is `40, then the ratio of daily wages
of a man and a boy will be.

Answer by LinnW(1048)

set b = boys rate in work projects per day
m = mens rate in work projects per day
10(2m + 4b) = 1
6(4m + 5b) = 1
Expanding the above
20m + 40b = 1
24m + 30b = 1
Since each equation equals 1
20m + 40b = 24m + 30b
add -20m to each side
40b = 24m -20m +30b
40b = 4m + 30b
add -30b to each side
10b = 4m
divide each side by 4
(10/4)b = m
(2.5)b = m
A mans output is 2.5 times a boys output
Try entering the below into www.wolframalpha.com
10(2m + 4b) = 1 ; 6(4m + 5b) = 1
The rate b = 1/90 and the rate m = 1/36 in terms of work projects per day.